In a given market, demand is described by the equation: QD=1800-10P and supply is described by: QS=200+10P.
(i) determine the equilibrium price and quantity
(ii) Graphically illustrate price and quantity
1
Expert's answer
2020-04-21T19:01:37-0400

"i) QD=QS;"

"1800-10P=200+10P;"

"1600=20P;\n\nP=" $ "80;" "Qd=1800-10\u00d780=" $ "1000;"

"Qs=200+10\u00d780=" $ "1000;"